A 100 kg block is pulled at a constant speed of 5 m/s across a horizontal force by an applied force of 122 N directed 37 o above the horizontal. The rate at which the force does work on the block is.


a. 810 Watts             

b. 490 Watts           

c. 530 Watts        

d. 704 Watt


1
Expert's answer
2022-02-28T10:43:09-0500

The power is given by

"P=\frac{W}{t}=\frac{Fd\cos\theta}{t}=Fv\cos\theta"

"P=\rm 122\: N*5.0\: m/s*\cos 37^\circ=487\: W"